OpenStudy (anonymous):

If y^3+x^(2)y-10=0, calculate y' (a,a)?

zepdrix (zepdrix):

\[\large\rm y^3+x^2y-10=0,\qquad\qquad y'(a,a)\]Implicit Differentiation, ya? :) What are you having trouble with? Looks like we'll need product rule for part of this.

OpenStudy (anonymous):

ohhh so it's 3y^(2)y' +2x+x^(2)y' =0?

OpenStudy (anonymous):

then what should i do?

zepdrix (zepdrix):

Ahh sorry ran off for a minute >.<

zepdrix (zepdrix):

Mmm your product rule looks a tiny bit off.

zepdrix (zepdrix):

\[\large\rm 3y^2y'+\color{red}{2x}+x^2y'=0\]

zepdrix (zepdrix):

\[\large\rm (x^2y)'=(x^2)'y+x^2(y)'\]Missing something in your middle term, ya? :)

OpenStudy (anonymous):

oh sorry i fell asleep, is the asnwer -1/2?